If |a+b|=|a| then find angle between |2a+b| and |b|.
 

Grade:9

1 Answers

Aditya Gupta
5 years ago
|a+b|=|a|
square both sides,
a^2+b^2+2a.b=a^2
or b.(b+2a)=0
here a.b means the dot product.
so the angle between b and 2a+b is pi/2.